OpenStudy (anonymous):

How do I write this in interval notation? 6 lxl < 1

OpenStudy (perl):

first I would solve the inequality for x

OpenStudy (p0sitr0n):

|x|< 1/6 since |x| > 0 then -1/6< x < 1/6 or x in (-1/6,1/6)

OpenStudy (anonymous):

\[ 6|x| < 1 \implies |x|<\frac 16 \]This leads to two equations: \[ x<\frac 16,\quad -x<\frac 16\implies x>-\frac 16 \]Combine them to get: \[ -\frac 16<x<\frac 16 \]The \(<\) becomes \((\), while \(\leq \) becomes \([\). The same logic applies to \(>\) and \(\geq\). Our interval is \((-1/6,1/6)\). We can say \(x\in(-1/6,1/6)\) to say that \(x\) is within the interval.